# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 4. Moment estimates for ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$
For $N\leq\widetilde{N}\in{\mathbbm{N}}$, consider the averaged partition function ${\mathcal{Z}}_{N}^{\beta_{\widetilde{N}}}(\varphi,\psi)$ as (2.14) with $\beta_{N}$ replaced by $\beta_{\widetilde{N}}=\beta_{\widetilde{N}}(\vartheta)$ satisfying (1.20). Fix $p,q\in(1+\infty)$ with $\frac{1}{p}+\frac{1}{q}=1$. Then there exist constants $C_{1},C_{2}\in(0,+\infty)$, such that for $\varphi,\psi\in C_{c}({\mathbbm{R}})$, and uniformly in $N\leq\widetilde{N}$,
(4.1) $\begin{array}[]{rl}\displaystyle\Big{|}{\mathbbm{E}}\Big{[}\Big{(}{\mathcal{Z}}^{\beta_{\widetilde{N}}}_{N}(\varphi,\psi)-{\mathbbm{E}}\big{[}{\mathcal{Z}}^{\beta_{\widetilde{N}}}_{N}(\varphi,\psi)\big{]}\Big{)}^{h}\Big{]}\Big{|}&\displaystyle\leq\frac{C_{1}}{\log(1+\frac{\widetilde{N}}{N})}\frac{1}{N^{\frac{h}{2}}}\lVert\varphi_{N}\rVert^{h}_{p}\lVert\psi_{N}\rVert^{h}_{\infty}\lVert\mathbbm{1}_{B_{N}}\rVert^{h}_{q}\\[5.0pt] &\displaystyle\leq\frac{C_{2}}{\log(1+\frac{\widetilde{N}}{N})}\lVert\varphi\rVert^{h}_{p}\lVert\psi\rVert^{h}_{\infty}\lVert\mathbbm{1}_{B}\rVert^{h}_{q},\end{array}$
@@
We will apply Theorem 4.1 in Section 7 with $h=4$.
@@
(4.2) $\begin{array}[]{l}\displaystyle\quad M_{N,\widetilde{N},h}^{\varphi,\psi}:={\mathbbm{E}}\Big{[}\Big{(}{\mathcal{Z}}_{N}^{\beta_{\widetilde{N}}}(\varphi,\psi)-{\mathbbm{E}}\big{[}{\mathcal{Z}}^{\beta_{\widetilde{N}}}_{N}(\varphi,\psi)\big{]}\Big{)}^{h}\Big{]}\\[10.0pt] \displaystyle=\frac{1}{N^{\frac{h}{2}}}{\mathbbm{E}}\Big{[}\Big{(}\sum_{r=1}^{\infty}\sum_{0<n_{1}<\ldots<n_{r}<N}q^{N}_{0,n_{1}}(\varphi,0)\zeta_{n_{1}}\Big{\{}\prod_{j=2}^{r}p_{n_{j-1},n_{j}}(0,0)\zeta_{n_{j}}\Big{\}}q^{N}_{n_{j},N}(0,\psi)\Big{)}^{h}\Big{]}.\end{array}$
Note that when expanding the $h$-fold product above, the summands are transition probabilities timing a sequence of disorders at the return times to 00 for $h$ random walks. By taking expectation, at each return time appearing in the summand, the disorder contributes a factor of ${\mathbbm{E}}[\zeta_{n}^{\#}]$, where ##\## is the number of random walks that return to 00 at time $n$. It is not hard to see that
(4.3) ${\mathbbm{E}}[\zeta_{n}]=0,\quad{\mathbbm{E}}[(\zeta_{n})^{2}]=\sigma_{\widetilde{N}}^{2}\sim\frac{2\pi}{\log\widetilde{N}},\quad\text{and}\quad|{\mathbbm{E}}[(\zeta_{n})^{l}]|\leq C\sigma_{\widetilde{N}}^{l}\quad\text{for}~{}l\geq 3.$
Hence, if a summand is non-zero, then at each return time, there must be at least two random walks returning to 00. Suppose that $l<m<n$ are three consecutive return times, and one of the $h$ random walks does not visit 00 at $m$. Its contribution is then only given by a transition probability ${\mathbf{P}}(S_{l}=x,S_{m}\neq 0,S_{n}=y)\leq p_{n-l}(y-x)$ supposing $S_{k}=x,S_{m}=z$. That is the reason why we only consider a pinning model associated to a random walk. For a general renewal process, we lack such a simple expression for transition kernels in the renewal decomposition.
@@
(4.4) ${\mathbbm{E}}[\zeta^{I}]:={\mathbbm{E}}[(\zeta_{n})^{|I(1)|}].$
(4.5) $Q_{t}(\boldsymbol{x},\boldsymbol{\widetilde{x}}):=\prod_{i=1}^{h}p_{t}(x_{i},\widetilde{x}_{i}),\quad Q_{t}^{N}(\varphi,\boldsymbol{x}):=\prod_{i=1}^{h}q_{0,t}^{N}(\varphi,x_{i}),\quad Q_{t}^{N}(\boldsymbol{x},\psi):=\prod_{i=1}^{h}q_{0,t}^{N}(x_{i},\psi).$
Also, for a given partition $I$, we denote
[…]
and for two consecutive partitions $I,J$, we denote
(4.7) $Q_{t}^{I,J}(\boldsymbol{x},\boldsymbol{\widetilde{x}}):=\mathbbm{1}_{\{\boldsymbol{x}\sim I,\boldsymbol{\widetilde{x}}\sim J\}}Q_{t}(\boldsymbol{x},\boldsymbol{\widetilde{x}}).$
@@
(4.8) $\begin{split}\quad M_{N,\widetilde{N},h}^{\varphi,\psi}=&\frac{1}{N^{\frac{h}{2}}}\sum\limits_{r=1}^{\infty}\sum\limits_{1\leq n_{1}<\cdots<n_{r}<n_{r+1}:=N\atop I_{1},\cdots,I_{r};\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{r}\in({\mathbbm{Z}})^{h}}Q_{n_{1}}^{N}(\varphi,\boldsymbol{x}_{1}){\mathbbm{E}}[\zeta^{I}]\\ &\times\prod_{i=2}^{r}Q_{n_{i}-n_{i-1}}^{I_{i-1},I_{i}}(\boldsymbol{x}_{i-1},\boldsymbol{x}_{i}){\mathbbm{E}}[\zeta^{I}]\times\mathbbm{1}_{\{\boldsymbol{x}_{r}\sim I_{r}\}}Q_{n_{r+1}-n_{r}}^{N}(\boldsymbol{x}_{r},\psi).\end{split}$
@@
(4.9) $|q_{0,N-n_{r}}^{N}(y,\psi)|\leq\|\psi\|_{\infty}\sum\limits_{v\in{\mathbbm{Z}}}p_{N-n_{r}}(v-y)\mathbbm{1}_{B}\Big{(}\frac{v}{\sqrt{N}}\Big{)}=\|\psi\|_{\infty}{\mathbf{P}}^{y}(S_{N-n_{r}}\in\sqrt{N}B).$
The probability is $O(1)$ if $N-n_{r}=O(N)$. Hence, there exists a constant $C$, such that uniformly for all $y$ and $1\leq n_{r}\leq N\leq n_{r+1}\leq 2N$,
$\|\psi\|_{\infty}{\mathbf{P}}^{y}(S_{N-n_{r}}\in\sqrt{N}B)\leq C\|\psi\|_{\infty}{\mathbf{P}}^{y}(S_{n_{r+1}-n_{r}}\in\sqrt{N}B)=C\|\psi\|_{\infty}q_{0,n_{r+1}-n_{r}}^{N}(y,\mathbbm{1}_{B_{N}}).$
@@
(4.10) $\begin{split}\Big{|}M_{N,\widetilde{N},h}^{\varphi,\psi}\Big{|}\leq&\frac{C^{h}\|\psi\|_{\infty}^{h}}{N^{\frac{h}{2}+1}}\sum\limits_{r=1}^{\infty}\sum\limits_{1\leq n_{1}<\cdots<n_{r}\leq n_{r+1}:=2N\atop I_{1},\cdots,I_{r};\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{r}\in({\mathbbm{Z}})^{h}}Q_{n_{1}}^{N}(\varphi,\boldsymbol{x}_{1}){\mathbbm{E}}[\zeta^{I_{1}}]\\ &\times\prod_{i=2}^{r}Q_{n_{i}-n_{i-1}}^{I_{i-1},I_{i}}(\boldsymbol{x}_{i-1},\boldsymbol{x}_{i}){\mathbbm{E}}[\zeta^{I_{i}}]\times\mathbbm{1}_{\{\boldsymbol{x}_{r}\sim I_{r}\}}Q_{n_{r+1}-n_{r}}^{N}(\boldsymbol{x}_{r},\mathbbm{1}_{B_{N}}).\end{split}$
We first single out a special case that some consecutive $I_{i},I_{i+1},\dots,I_{j}$ are the same with $|I_{i}|=\cdots=|I_{j}|=h-1$. Note that in this case there exist $1\leq m\neq l\leq h$, such that, $I_{k}(1)=\{m,l\}$ for $k=i,i+1,\cdots,j$, namely, random walks with indices $m$ and $l$ keep returning to 00 consecutively at all return times, while other walks never return to 00.
@@
(4.11) $U^{I}_{t-s,\widetilde{N}}(\boldsymbol{x},\boldsymbol{\widetilde{x}}):=\mathbbm{1}_{\{\boldsymbol{x},\boldsymbol{\widetilde{x}}\sim I\}}\sum\limits_{k=1}^{\infty}{\mathbbm{E}}[\zeta^{2}]^{k}\sum\limits_{n_{0}:=s<n_{1}<\cdots<n_{k}:=t\atop\boldsymbol{y}_{i}\in({\mathbbm{Z}})^{h},\boldsymbol{y}_{0}:=\boldsymbol{x},\boldsymbol{y}_{k}:=\boldsymbol{\widetilde{x}}}\prod\limits_{i=1}^{k}Q_{n_{i}-n_{i-1}}^{I,I}(\boldsymbol{y}_{i-1},\boldsymbol{y}_{i}).$
Note that here $\zeta$ depends on $\widetilde{N}$ and $U_{0,\widetilde{N}}^{I}(\boldsymbol{x},\boldsymbol{\widetilde{x}})$ is understood as $\mathbbm{1}_{\{\boldsymbol{x},\boldsymbol{\widetilde{x}}\sim I\}}$. Then in (4.10), for any given first and last return times $s,t$, we can contract all consecutive $I_{i}=I_{i+1}=\cdots=I_{j}:=I$ for any $j-i\in{\mathbbm{N}}$ to a single kernel $U_{t-s}^{I}(\cdot,\cdot)$. Recall $\overline{U}_{N}(n)$ from (2.24), and we write
(4.12) $U_{N}(n)=\overline{U}_{N}(n)/\sigma_{N}^{2}.$
@@
(4.13) $U^{I}_{t-s,\widetilde{N}}(\boldsymbol{x},\boldsymbol{\widetilde{x}})\leq U_{\widetilde{N}}(t-s)\prod_{i\in\{1,\cdots,h\}\backslash\{m,l\}}p_{s,t}(x_{i},\widetilde{x}_{i}).$
@@
(4.14) $Q_{\lambda,N}^{I,J}(\boldsymbol{x},\boldsymbol{\widetilde{x}})$ $:=\sum\limits_{n=1}^{2N}e^{-\lambda n}Q_{n}^{I,J}(\boldsymbol{x},\boldsymbol{\widetilde{x}}),\quad\boldsymbol{x},\boldsymbol{\widetilde{x}}\in({\mathbbm{Z}})^{h},$
(4.15) $U_{\lambda,N,\widetilde{N}}^{J}(\boldsymbol{x},\boldsymbol{\widetilde{x}})$ $:=\sum\limits_{n=1}^{2N}e^{-\lambda n}U_{n,\widetilde{N}}^{J}(\boldsymbol{x},\boldsymbol{\widetilde{x}}),\quad\boldsymbol{x},\boldsymbol{\widetilde{x}}\in({\mathbbm{Z}})^{h}.$
Then we define the operators from $({\mathbbm{Z}})^{h}\rightarrow({\mathbbm{Z}})^{h}$ by
(4.16) $P_{\lambda}^{I,J}=P_{\lambda,N,\widetilde{N}}^{I,J}:=\begin{cases}Q_{\lambda,N}^{I,J},\quad&\text{if}~{}|J|<h-1,\\ Q_{\lambda,N}^{I,J}U_{\lambda,N,\widetilde{N}}^{J},\quad&\text{if}~{}|J|=h-1.\end{cases}$
Now in each summand on the right-hand side of (4.10), we insert a term $e^{2\lambda N}e^{-\lambda\sum_{i=1}^{r+1}(n_{i}-n_{i-1})}\geq 1$, where $\lambda>0$ will be determined later. For the initial and terminal times, there is no constraint on the positions of the random walks, and we denote the “partition” by ***. Finally, (4.10) can be further bounded above by
(4.17) $\Big{|}M_{N,\widetilde{N},h}^{\varphi,\psi}\Big{|}\leq\frac{C_{h}\|\psi\|_{\infty}^{h}e^{2\lambda N}}{N^{\frac{h}{2}+1}}\sum\limits_{r=1}^{\infty}\sum_{I_{1},\cdots,I_{r}}\big{\langle}\varphi_{N}^{\otimes h},P^{*,I_{1}}_{\lambda}P^{I_{1},I_{2}}_{\lambda}\cdots P_{\lambda}^{I_{r-1},I_{r}}Q_{\lambda}^{I_{r},*}\mathbbm{1}_{B_{N}}^{\otimes h}\big{\rangle}\prod_{i=1}^{r}{\mathbbm{E}}[\zeta^{I_{i}}].$
@@
(4.18) $\begin{split}\Big{|}M_{N,\widetilde{N},h}^{\varphi,\psi}\Big{|}\leq\frac{C_{h}\|\psi\|_{\infty}^{h}\|\varphi\|_{p}^{h}e^{2\lambda N}}{N^{\frac{h}{2}+1}}\sum\limits_{r=1}^{\infty}&\sum_{I_{1},\cdots,I_{r}}\Big{\|}P_{\lambda}^{*,I_{1}}\Big{\|}_{\ell^{q}\to\ell^{q}}\Big{\|}P_{\lambda}^{I_{1},I_{2}}\Big{\|}_{\ell^{q}\to\ell^{q}}\cdots\\ &\cdots\Big{\|}P_{\lambda}^{I_{r-1},I_{r}}\Big{\|}_{\ell^{q}\to\ell^{q}}\Big{\|}Q_{\lambda}^{I_{r},*}\Big{\|}_{\ell^{q}\to\ell^{q}}\Big{\|}\mathbbm{1}_{B_{N}}^{\otimes h}\Big{\|}_{q}\prod\limits_{i=1}^{r}{\mathbbm{E}}[\zeta^{I_{i}}],\end{split}$
where the norm $\|\cdot\|_{\ell^{q}\to\ell^{q}}$ for the operator $A:({\mathbbm{Z}})^{h}\rightarrow({\mathbbm{Z}})^{h}$ is defined by

### v3

@@ 4 Moment estimates for ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$
For $N\leq\widetilde{N}\in{\mathbbm{N}}$, consider ${\mathcal{Z}}_{N}^{\beta_{\widetilde{N}}}(\varphi,\psi)$ as (2.14) with $\beta_{N}$ replaced by $\beta_{\widetilde{N}}=\beta_{\widetilde{N}}(\vartheta)$ satisfying (1.20). Fix $p,q\in(1+\infty)$ with $\frac{1}{p}+\frac{1}{q}=1$ and $h\geq 3$. Then there exist constants $C_{1},C_{2}\in(0,+\infty)$ depending on $p,q,h$ and $\omega$, such that for $\varphi,\psi\in C_{c}({\mathbbm{R}})$, and uniformly in $N\leq\widetilde{N}$,
$\begin{array}[]{rl}\displaystyle\Big|{\mathbbm{E}}\Big[\Big({\mathcal{Z}}^{\beta_{\widetilde{N}}}_{N}(\varphi,\psi)-{\mathbbm{E}}\big[{\mathcal{Z}}^{\beta_{\widetilde{N}}}_{N}(\varphi,\psi)\big]\Big)^{h}\Big]\Big|&\displaystyle\leq\frac{C_{1}}{\log(1+\frac{\widetilde{N}}{N})}\frac{1}{N^{\frac{h}{2}}}\lVert\varphi_{N}\rVert^{h}_{p}\lVert\psi_{N}\rVert^{h}_{\infty}\lVert\mathbbm{1}_{B_{N}}\rVert^{h}_{q}\\[5.0pt] &\displaystyle\leq\frac{C_{2}}{\log(1+\frac{\widetilde{N}}{N})}\lVert\varphi\rVert^{h}_{p}\lVert\psi\rVert^{h}_{\infty}\lVert\mathbbm{1}_{B}\rVert^{h}_{q},\end{array}$ (4.1)
@@
We will use Theorem 4.1 with $h=4$ to prove Lemma 6.2. Therein, the coarse-grained disorder can be viewed as ${\mathcal{Z}}_{N}^{\beta_{\widetilde{N}}}(\varphi,\psi)$ with $N\in\{\varepsilon\widetilde{N},2\varepsilon\widetilde{N}\}$, where $\varepsilon>0$ is small and will be sent to 0, leading to $N\leq\widetilde{N}$ in Theorem 4.1. Moreover, we can see from the proof that for compactly supported $\varphi,\psi$, the assumption of continuity can be weakened to boundedness and measurability.
@@
$\begin{array}[]{l}\displaystyle\quad M_{N,\widetilde{N},h}^{\varphi,\psi}:={\mathbbm{E}}\Big[\Big({\mathcal{Z}}_{N}^{\beta_{\widetilde{N}}}(\varphi,\psi)-{\mathbbm{E}}\big[{\mathcal{Z}}^{\beta_{\widetilde{N}}}_{N}(\varphi,\psi)\big]\Big)^{h}\Big]\\[10.0pt] \displaystyle=\frac{1}{N^{\frac{h}{2}}}{\mathbbm{E}}\Big[\Big(\sum_{r=1}^{\infty}\sum_{0<n_{1}<\ldots<n_{r}<N}q^{N}_{0,n_{1}}(\varphi,0)\zeta_{n_{1}}\Big\{\prod_{j=2}^{r}p_{n_{j-1},n_{j}}(0,0)\zeta_{n_{j}}\Big\}q^{N}_{n_{j},N}(0,\psi)\Big)^{h}\Big].\end{array}$ (4.2)
Note that when expanding the $h$-fold product above, the summands are transition probabilities timing a sequence of disorders at the return times to 0 for $h$ random walks. By taking expectation, at each return time appearing in the summand, the disorder contributes a factor of ${\mathbbm{E}}[\zeta_{n}^{\#}]$, where #\# is the number of random walks that return to 0 at time $n$. We have that (see equation (6.7) in [11])
${\mathbbm{E}}[\zeta_{n}]=0,\quad{\mathbbm{E}}[(\zeta_{n})^{2}]=\sigma_{\widetilde{N}}^{2}\sim\frac{2\pi}{\log\widetilde{N}},\quad\text{and}\quad|{\mathbbm{E}}[(\zeta_{n})^{l}]|\leq C_{h}\sigma_{\widetilde{N}}^{l}\quad\text{for}~3\leq l\leq h.$ (4.3)
Hence, if a summand is non-zero, then at each return time, there must be at least two random walks returning to 0. Suppose that $l<m<n$ are three consecutive return times, and one of the $h$ random walks does not visit 0 at $m$. Its contribution is then only given by a transition probability ${\mathbf{P}}(S_{l}=x,S_{m}\neq 0,S_{n}=y)\leq p_{n-l}(y-x)$ supposing $S_{l}=x,S_{n}=y$. That is the reason why we only consider a pinning model associated to a random walk. For a general renewal process, we lack such a simple expression for transition kernels in the renewal decomposition.
At a return time $n$, the $h$ random walks can visit $m$ positions, where $1\leq m\leq h-1$. Note that $m$ cannot be $h$, since otherwise we have ${\mathbbm{E}}[(\zeta_{n})^{1}]=0$. We then partition the indices $\{1,\cdots,h\}$ into $m$ sets $I:=\{I(k):k=1,\cdots,m\}$, such that $\cup_{k=1}^{m}I(k)=\{1,\cdots,h\}$ and $I(j)\cap I(k)=\emptyset,\forall j\neq k$. For $i,j\in\{1,\cdots,h\}$, we write $i\overset{I}{\sim}j$ if $i,j\in I(k)$ for some $1\leq k\leq m$. In particular, we use $I(1)$ to record the indices of random walks that visit 0. Apparently, we only need to consider the case $|I(1)|\geq 2$. Hence, we have that at return time $n$,
${\mathbbm{E}}[\zeta^{I}]:={\mathbbm{E}}[(\zeta_{n})^{|I(1)|}].$ (4.4)
$Q_{t}(\boldsymbol{x},\boldsymbol{\widetilde{x}}):=\prod_{i=1}^{h}p_{t}(x_{i},\widetilde{x}_{i}),\quad Q_{t}^{N}(\varphi,\boldsymbol{x}):=\prod_{i=1}^{h}q_{0,t}^{N}(\varphi,x_{i}),\quad Q_{t}^{N}(\boldsymbol{x},\psi):=\prod_{i=1}^{h}q_{0,t}^{N}(x_{i},\psi).$ (4.5)
Also, for a given partition $I$, we denote
$\boldsymbol{x}\sim I,\quad\text{if}~x_{j}=x_{k}\Longleftrightarrow j\overset{I}{\sim}k,\quad\forall j,k=1,\cdots,h,$ (4.6)
and for two consecutive partitions $I,J$, we denote
@@
$\begin{split}\quad M_{N,\widetilde{N},h}^{\varphi,\psi}=&\frac{1}{N^{\frac{h}{2}}}\sum\limits_{r=1}^{\infty}\sum\limits_{1\leq n_{1}<\cdots<n_{r}<n_{r+1}:=N\atop I_{1},\cdots,I_{r};\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{r}\in({\mathbbm{Z}})^{h}}Q_{n_{1}}^{N}(\varphi,\boldsymbol{x}_{1}){\mathbbm{E}}[\zeta^{I}]\\ &\times\prod_{i=2}^{r}Q_{n_{i}-n_{i-1}}^{I_{i-1},I_{i}}(\boldsymbol{x}_{i-1},\boldsymbol{x}_{i}){\mathbbm{E}}[\zeta^{I}]\times\mathbbm{1}_{\{\boldsymbol{x}_{r}\sim I_{r}\}}Q_{n_{r+1}-n_{r}}^{N}(\boldsymbol{x}_{r},\psi).\end{split}$ (4.8)
@@
$|q_{0,N-n_{r}}^{N}(y,\psi)|\leq\|\psi\|_{\infty}\sum\limits_{v\in{\mathbbm{Z}}}p_{N-n_{r}}(v-y)\mathbbm{1}_{B}\Big(\frac{v}{\sqrt{N}}\Big)=\|\psi\|_{\infty}{\mathbf{P}}^{y}(S_{N-n_{r}}\in\sqrt{N}B).$ (4.9)
The probability at the right-hand side above is of $O(1)$ if $N-n_{r}=O(N)$. Hence, there exists a constant $C$, such that uniformly for all $y$ and $1\leq n_{r}\leq N\leq n_{r+1}\leq 2N$,
${\mathbf{P}}^{y}(S_{N-n_{r}}\in\sqrt{N}B)\leq C{\mathbf{P}}^{y}(S_{n_{r+1}-n_{r}}\in\sqrt{N}B)=Cq_{0,n_{r+1}-n_{r}}^{N}(y,\mathbbm{1}_{B_{N}}).$
@@
$\begin{split}\Big|M_{N,\widetilde{N},h}^{\varphi,\psi}\Big|\leq&\frac{C^{h}\|\psi\|_{\infty}^{h}}{N^{\frac{h}{2}+1}}\sum\limits_{r=1}^{\infty}\sum\limits_{1\leq n_{1}<\cdots<n_{r}\leq n_{r+1}:=2N\atop I_{1},\cdots,I_{r};\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{r}\in({\mathbbm{Z}})^{h}}Q_{n_{1}}^{N}(\varphi,\boldsymbol{x}_{1}){\mathbbm{E}}[\zeta^{I_{1}}]\\ &\times\prod_{i=2}^{r}Q_{n_{i}-n_{i-1}}^{I_{i-1},I_{i}}(\boldsymbol{x}_{i-1},\boldsymbol{x}_{i}){\mathbbm{E}}[\zeta^{I_{i}}]\times\mathbbm{1}_{\{\boldsymbol{x}_{r}\sim I_{r}\}}Q_{n_{r+1}-n_{r}}^{N}(\boldsymbol{x}_{r},\mathbbm{1}_{B_{N}}).\end{split}$ (4.10)
We first single out a special case that some consecutive $I_{i},I_{i+1},\dots,I_{j}$ are the same with $|I_{i}|=\cdots=|I_{j}|=h-1$. Note that in this case there exist $1\leq m\neq l\leq h$, such that, $I_{k}(1)=\{m,l\}$ for all $k=i,i+1,\cdots,j$, namely, random walks with indices $m$ and $l$ keep returning to 0 consecutively at all return times, while other walks never return to 0.
@@
$U^{I}_{t-s,\widetilde{N}}(\boldsymbol{x},\boldsymbol{\widetilde{x}}):=\mathbbm{1}_{\{\boldsymbol{x},\boldsymbol{\widetilde{x}}\sim I\}}\sum\limits_{k=1}^{\infty}{\mathbbm{E}}[\zeta^{2}]^{k}\sum\limits_{n_{0}:=s<n_{1}<\cdots<n_{k}:=t\atop\boldsymbol{y}_{i}\in({\mathbbm{Z}})^{h},\boldsymbol{y}_{0}:=\boldsymbol{x},\boldsymbol{y}_{k}:=\boldsymbol{\widetilde{x}}}\prod\limits_{i=1}^{k}Q_{n_{i}-n_{i-1}}^{I,I}(\boldsymbol{y}_{i-1},\boldsymbol{y}_{i}).$ (4.11)
[…]
$U_{N}(n)=\overline{U}_{N}(n)/\sigma_{N}^{2}.$ (4.12)
@@
$U^{I}_{t-s,\widetilde{N}}(\boldsymbol{x},\boldsymbol{\widetilde{x}})\leq U_{\widetilde{N}}(t-s)\prod_{i\in\{1,\cdots,h\}\backslash\{m,l\}}p_{s,t}(x_{i},\widetilde{x}_{i}).$ (4.13)
@@
[…]
Then we define the operators from $({\mathbbm{Z}})^{h}$ to $({\mathbbm{Z}})^{h}$ by
$P_{\lambda}^{I,J}=P_{\lambda,N,\widetilde{N}}^{I,J}:=\begin{cases}Q_{\lambda,N}^{I,J},\quad&\text{if}~|J|<h-1,\\ Q_{\lambda,N}^{I,J}U_{\lambda,N,\widetilde{N}}^{J},\quad&\text{if}~|J|=h-1.\end{cases}$ (4.16)
Now in each summand on the right-hand side of (4.10), we insert a term $e^{-\lambda\sum_{i=1}^{r+1}(n_{i}-n_{i-1})}$
$\times e^{2\lambda N}\geq 1$, where $\lambda>0$ will be determined later. For the initial and terminal times, there is no constraint on the positions of the random walks, and we denote the “partition” by *. Finally, (4.10) can be further bounded above by
$\Big|M_{N,\widetilde{N},h}^{\varphi,\psi}\Big|\leq\frac{C_{h}\|\psi\|_{\infty}^{h}e^{2\lambda N}}{N^{\frac{h}{2}+1}}\sum\limits_{r=1}^{\infty}\sum_{I_{1},\cdots,I_{r}}\big\langle\varphi_{N}^{\otimes h},P^{*,I_{1}}_{\lambda}P^{I_{1},I_{2}}_{\lambda}\cdots P_{\lambda}^{I_{r-1},I_{r}}Q_{\lambda}^{I_{r},*}\mathbbm{1}_{B_{N}}^{\otimes h}\big\rangle\prod_{i=1}^{r}{\mathbbm{E}}[\zeta^{I_{i}}].$ (4.17)
Here the summation is over all partitions $I_{1},\cdots,I_{r}$ of $\{1,\cdots,h\}$, such that $|I_{i}|\leq h-1$ and no consecutive $I_{i}=I_{i+1}$ with $|I_{i}|=h-1$. The operation $\big\langle\cdot,\cdot\big\rangle$ is the classic notation for integral on $({\mathbbm{Z}})^{h}\otimes({\mathbbm{Z}})^{h}$. Given a function $f:{\mathbbm{Z}}\rightarrow{\mathbbm{R}}$, for $\boldsymbol{x}\in({\mathbbm{Z}})^{h}$, $f^{\otimes h}(\boldsymbol{x}):=\prod_{i=1}^{h}f(x_{i})$.
@@
$\begin{split}\Big|M_{N,\widetilde{N},h}^{\varphi,\psi}\Big|\leq\frac{C_{h}\|\psi\|_{\infty}^{h}\|\varphi\|_{p}^{h}e^{2\lambda N}}{N^{\frac{h}{2}+1}}&\sum\limits_{r=1}^{\infty}\sum_{I_{1},\cdots,I_{r}}\Big\|P_{\lambda}^{*,I_{1}}\Big\|_{\ell^{q}\to\ell^{q}}\Big\|P_{\lambda}^{I_{1},I_{2}}\Big\|_{\ell^{q}\to\ell^{q}}\cdots\\ \cdots&\Big\|P_{\lambda}^{I_{r-1},I_{r}}\Big\|_{\ell^{q}\to\ell^{q}}\Big\|Q_{\lambda}^{I_{r},*}\Big\|_{\ell^{q}\to\ell^{q}}\Big\|\mathbbm{1}_{B_{N}}^{\otimes h}\Big\|_{q}\prod\limits_{i=1}^{r}{\mathbbm{E}}[\zeta^{I_{i}}],\end{split}$ (4.18)
where the norm $\|\cdot\|_{\ell^{q}\to\ell^{q}}$ for an operator $A:({\mathbbm{Z}})^{h}\rightarrow({\mathbbm{Z}})^{h}$ is defined by